OpenStudy (anonymous):

@Data_LG2 @ganeshie8 Do you know how to do this problem?

ganeshie8 (ganeshie8):

compare it with the standard form of equation of circle : \[\large (x-h)^2 + (y-k)^2 = r^2\] center = \(\large (h,k)\) radius = \(\large r\)

OpenStudy (anonymous):

so since the formula is already filled out in the problem, I basically plug in the answer choices given into x and y- until you get 9/4 , correct?

ganeshie8 (ganeshie8):

nope, we just compare both the equations : apples to apples and coconuts to coconutes

ganeshie8 (ganeshie8):

Your given equation : \[\large (x-1)^2 + \left(y + \frac{3}{2}\right)^2 = \frac{9}{4}\]

ganeshie8 (ganeshie8):

can we rewrite like this : \[\large (x-1)^2 + \left(y -(- \frac{3}{2})\right)^2 = \frac{9}{4}\] ?

ganeshie8 (ganeshie8):

Next we compare it with below standard equation : \[\large (x-h)^2 + (y-k)^2 = r^2 \]

ganeshie8 (ganeshie8):

You get : h = 1 k = -3/2 r^2 = 9/4

ganeshie8 (ganeshie8):

So, center = \(\large (h, k) = \left(1, -\frac{3}{2}\right)\)

ganeshie8 (ganeshie8):

find the raidus

OpenStudy (anonymous):

@ganeshie8 Sorry for the delay, I stepped away for a couple minutes and then OS went to updates when I wanted to comment. Thank you for partially working it out, thus guiding me on the proper direction to take in order to solve for radius, which is \[\frac{ 3 }{ 2 } (:\]